OpenStudy (anonymous):

what should you substitute for x in the second equation (bottom equation) in order to solve the system by the substitution method? write your answer as a polynomial in descending order of terms. 2x -1/3 y = -2 3x + y= -1 (slash is a fraction bar)

OpenStudy (campbell_st):

2x - 1/3 y = -2 make x the subject 2x = 1/3 y - 2 x = 1/2(1/3y - 2) or x= 1/6 y - 1 when substituted into (2) you would have 3(1/6 y - 1) +y = -1 or 3/2y - 3 = -1 hope this helps... the numbers aren't very friendly